OpenStudy (anonymous):

Which of the following shows the correct steps to find the value of 16 to the power of 1 over 4?

OpenStudy (anonymous):

16 ^1/4=2^4)1/4= (2) 4 1/4= 2 16^1/4= (4^4)1/4= (4) 4 1/4= 4 16^1/4= (2^8)1/4= (8) 4 1/4= 4 16^1/4= (8^2)1/4= (2) 2 1/4=8